A boy in a hot air balloon drops his cellular phone when the balloon is 150 ft above the ground and rising at the rate of 10 ft/s.

a.How long will it take the cellular phone to reach the ground?

b. With what speed will it strike the ground?


1
Expert's answer
2022-04-19T04:40:55-0400

a. "h(t)=150+10t-16t^2."

"h(t)=0 \to 150+10t-16t^2=0 \to t=\frac{5+\sqrt{25+9600}}{32}=3.39 s."

The cellular reaches the ground in 3.39 seconds.


b. "v(t)=10-32t."

"v(3.39)=10-32*3.39=-98.48ft\/s."